Please help I only have 10 min to answer


media%2Fe59%2Fe592cffd-9c46-4d06-82c1-df
Answer #1

A.) Total number of marbles = 8+9+10 = 27

Probability of drawing a red marble first = 9/27 (since there are 9 red marbles and we can draw any one of those 9 out of the total 27)

Next, since that marble is put in the bag again since we have replacement here, probability of drawing a green marble now is 10/27

And: So required probability = 9*10/27*27 = 0.123457

B.) Here the marbles are drawn from the bag without replacement. Probability of first drawing a yellow marble = 8/27

Next, probability of drawing a white marble = 9/26 (denominator is now 26 since one yellow marble is already drawn)

Ans: so the required probability = 8*9/27*26 = 0.102564